OpenStudy (anonymous):

Use the product or quotient properties to write each equation in simplified form: 3sqrt{9p^7q^5}

OpenStudy (anonymous):

\[3\sqrt{9p^7q^5}\]

OpenStudy (anonymous):

that looks better

OpenStudy (anonymous):

\[3\sqrt{9p^7q^5}\] \[3\sqrt{3^2p^6q^4*pq}\] \[9p^3q^2\sqrt{pq}\] In step 2, I separated it into numbers that could be square rooted so I could take it out of the equation.

OpenStudy (anonymous):

because the 7 and 5 can not be squared?

OpenStudy (anonymous):

Yes. Find the lowest number that can be divided by 2 (because exponent laws say that division means subtraction of exponents and 2 is the index of the radical). In this case, 6 and 4.

OpenStudy (anonymous):

*7 and 5 cannot be square rooted to give an integer result.

OpenStudy (anonymous):

is it that way for any one that can not be squared?

OpenStudy (anonymous):

Yes. Remember for a square root\(\sqrt{blah}\), the index is 2. (That's the little number. So it would actually be \[\sqrt[2]{blah}\] Now we want to square root variables to take them out. You do this by dividing them (technically, it's because square roots can be written as exponents). \(x^6\) can be divided by 2 to give \(x^3\) (Notice the exponent is being divided). \(x^7\) cannot be, so split it up to \(x^6 * x\). You can now take out the \(x^3\) but NOT the \(x\) so leave that inside.

OpenStudy (anonymous):

these are so confusing to me, I have a bunch to figure out and a test on it tuesday at noon

OpenStudy (anonymous):

If you're willing to learn, I highly recommend the khan academy videos. Specifically: http://www.khanacademy.org/math/algebra/exponents-radicals/v/simplifying-radical-expressions1 http://www.khanacademy.org/math/algebra/exponents-radicals/v/simplifying-radical-expressions-2 http://www.khanacademy.org/math/algebra/exponents-radicals/v/simplifying-radical-expressions-3 Together, it will take you probably about 20-30 minutes to watch, so it's worth a look if you some time tonight.